# Optimal Protocols for Continual Learning via Statistical Physics and Control Theory

### v1

@@ 3.1 Experiments on synthetic data
We use the equations derived in Sec. 2 to study optimal replay during the second phase of training. To this end, we take the task-selection variable as our control $u(\alpha)\!=\!t_{c}(\alpha)\!\in\!\{1,2\}$, while we set $t_{c}\!=\!1$ during the first training phase. The result of the optimisation in Eq. 9 strikes the balance between training on the new task and replaying the old task. We do not enforce any constraints on the number of samples from task $1$ to use in the second phase, therefore our method provides both the optimal fraction of replayed samples and the optimal task ordering, depending on the time window $\alpha_{F}$. Fig. 2 compares the learning dynamics of three different strategies, depicting the loss on task $1$ (full orange line), task $2$ (dashed green line), and their average (dotted black line) as a function of the training time $\alpha$. Numerical simulations—marked by cross symbols—are in excellent agreement with our theory, with deviations smaller than $1/\sqrt{N}$ and hence compatible with finite-size effects.
@@
To understand the structure of the optimal strategy, we examine its performance in relation to task similarity $\gamma$. Fig. 3a) depicts the average loss at the end of training as a function of $\gamma$. For the no-replay strategy, we reproduce the findings from previous works (Lee et al., 2021; 2022): the highest error occurs at intermediate task similarity. Lee et al. (2022) explained this non-monotonicity as a trade-off between node re-use and node activation. Specifically, for small $\gamma$, there is minimal interference between tasks, and one hidden neuron predominantly aligns with the new task, while the other neuron retains the knowledge of the old task, leading to task specialisation. At large $\gamma$, features from task $1$ are reused for task $2$, avoiding forgetting. However, at intermediate $\gamma$, interference is maximal, and both neurons quickly align with task $2$, while task $1$ is forgotten. Remarkably, Fig. 3a) shows that replay reverses this trend: the minimal error occurs at intermediate $\gamma$. To explain this nontrivial behaviour, we must first understand the optimal replay protocol.
@@
The optimal replay dynamics is illustrated in panels b-e) of Fig. 3 for different values of $\gamma$, displaying a highly structured protocol. We can interpret this structure a posteriori: an initial focus phase without replay is followed by a revision phase involving interleaved replay. The transition between these two phases corresponds approximately to the point at which the loss on the new task matches the loss on the old one. To investigate the significance of this structure, we also test an interleaved strategy, plotted in Fig. 3a), where the task ordering in the second training phase is fully randomised while maintaining the same overall replay fraction of the optimal strategy. This protocol has a performance gap compared to the optimal one, showing the importance of a properly structured replay scheme. Additionally, we test a “pseudo-optimal” variant, where the focus phase is retained, but the revision phase is randomised. This variant performs comparably to the optimal strategy, suggesting that while the specific order of the revision phase is largely unimportant, it is key to precede it with a training phase on the new task.
@@
We can now attempt to understand the inverted non-monotonic behaviour of the average loss as a function of $\gamma$ under the optimal protocol. First, as shown in Fig. 8 of Appendix C, the optimal protocol achieves a good level of node specialisation across all values of $\gamma$. Thus, replay prevents the task interference that typically causes performance deterioration at intermediate $\gamma$. The non-monotonic behaviour of the optimal curve in Fig. 3a) arises from a different origin, involving two opposing effects related to the first-layer weights and the readout. The initial decrease of the loss with $\gamma$ is quite intuitive, as only minimal knowledge can be transferred from task $1$ to task $2$ when $\gamma$ is small. Consequently, the focus phase on task $2$ must be longer for smaller $\gamma$, leaving less time to revise task $1$, thereby reducing performance. On the other hand, the performance decrease observed in Fig. 3a) for $\gamma>0.3$ is more subtle and is related to the readout layer. Once the two hidden neurons have specialised—each aligning with one of the teacher vectors—we expect the readout weights corresponding to the incorrect teacher to be suppressed. Specifically, if $\bm{w}_{1}=\bm{w}_{*}^{(1)}$ and $\bm{w}_{2}=\bm{w}_{*}^{(2)}$, the learning dynamics should drive the readout weights $\bm{v}^{(1)}=(v^{(1)}_{1},v^{(1)}_{2})^{\top}$ and $\bm{v}^{(2)}=(v^{(2)}_{1},v^{(2)}_{2})^{\top}$ towards $\bm{v}^{(1)}=(1,0)^{\top}$ and $\bm{v}^{(2)}=(0,1)^{\top}$, representing full recovery of the teacher network. As shown in Fig. 8 of Appendix C, the time required to suppress the off-diagonal weights $v^{(1)}_{2}$ and $v^{(2)}_{1}$ increases as $\gamma\to 1$. This is intuitive, as higher task similarity $\gamma$ reduces the distinction between tasks, slowing the suppression of the off-diagonal weights. In Appendix A, we derive analytically the convergence timescale $\alpha_{\rm conv}$ of the readout layer, showing that
@@
While the above results were obtained with constant learning rate, annealing schedules are widely used in machine learning. Thus, it is relevant to study the optimal interplay between replay and learning rate scheduling. Optimal learning rate dynamics have been studied with a similar approach in Saad & Rattray (1997). We jointly optimise the task-selection protocol together with the learning rate to investigate its impact on continual learning. Fig. 4 shows the optimal learning rate schedule for task similarity $\gamma\!=\!0.3$ in the second training phase of duration $\alpha_{F}\!=\!25$. Similarly as for constant learning rate (panel c) of Fig. 3), optimal task-selection is characterised by an initial focus phase. Notably, this phase coincides with a strong annealing of the learning rate to achieve optimal performance. Intuitively, when learning the new task, the learning rate starts high and is gradually decreased over time. Interestingly, while entering the revision phase, the optimal learning rate schedule exhibits a highly nontrivial structure (see Fig. 4). Indeed, despite the optimal learning rate curve is unique, we find that effectively it can be seen as two different curves, associated to the respective tasks. In practice, the optimal learning rate curve “jumps” between these two curves according to the task selected at a given training time. Overall, the joint optimisation over task selection and learning rate provides a significant improvement in performance. Indeed, the average loss obtained by this strategy—shown in Fig. 9 of Appendix C—exhibits a large gap compared to all the other protocols.
@@
We also consider a multi-task setting (see Fig. 5) where both tasks must be learned from scratch within a fixed number of epochs corresponding to a total training time $\alpha_{F}$. We first consider sequential learning, i.e., training only on task $1$ for $\alpha<\alpha_{F}/2$, then only on task $2$, or vice versa. As shown in Fig. 5a) sequential learning leads to catastrophic forgetting, with the worst performance observed at intermediate task similarity. In contrast, using a randomly interleaved strategy, where examples from both tasks are presented in equal proportion but in random order, shows significant improvement. This approach can exploit task similarity, leading to a monotonic decrease in average loss as $\gamma$ increases. The optimal strategy, displayed in Fig. 5b-e) for various values of $\gamma$, follows a structured interleaved protocol that further enhances performance. Contrary to the continual learning framework, the optimal structure gives only marginal gain over the plain interleaved stratetegy and does not offer a clear a posteriori interpretation.
@@ 3.2 Experiments on real data
We consider the experimental framework established in Ramasesh et al. (2020); Lee et al. (2022) for the study of task similarity in relation to catastrophic forgetting. We use the Fashion-MNIST dataset (Xiao et al., 2017) to generate upstream and downstream tasks: The upstream dataset—$\mathcal{D}_{1}=\{\boldsymbol{x}_{i}^{(1)},y_{i}^{(1)}\}_{i}$—consists in a pair of classes from the standard dataset, while the downstream dataset is generated by a linear interpolation of the upstream dataset with a second auxiliary dataset—$\tilde{\mathcal{D}}=\{\tilde{\boldsymbol{x}}_{i},\tilde{y}_{i}\}_{i}$—containing a new pair of classes,
$\mathcal{D}_{2}=\{\boldsymbol{x}_{i}^{(2)},y_{i}^{(2)}\}_{i}=\{\gamma\boldsymbol{x}_{i}^{(1)}+(1-\gamma)\tilde{\boldsymbol{x}}_{i},\gamma y_{i}^{(1)}+(1-\gamma)\tilde{y}_{i}\}_{i}$ (11)
@@
where the parameter $\gamma$ controls the task similarity. We then train a standard two-layer feedforward ReLU neural network on the two datasets using online SGD on a squared error loss. We consider a dynamical architecture (Zhou et al., 2012; Rusu et al., 2016) where the readout weights are changed switching from one task to another, but the hidden layer is shared. During training, we apply the three strategies discussed in the previous sections: a no-replay strategy, a strategy with interleaved replay, and a “pseudo-optimal” strategy. Recall that the latter is inspired by the optimal protocol derived in the previous section. It consists of an initial phase of training exclusively on the new task until performance on both tasks becomes comparable, followed by a phase of interleaved replay. Crucially, this protocol can be easily implemented in practice, as it only requires an estimate of the generalization error on the two tasks, which can be obtained in real-world settings.
@@
Figure 6: Training dynamics. Training curves on the modified fashion MNIST task at similarity $\gamma=0.5$. The network is trained for $10.000$ epochs on the first task before switching to the second task and being trained for additional $10.000$ epochs. The results are obtained from $100$ realisations of the problem. The first three panels show the test loss on task $1$ (solid orange), task $2$ (dashed green), and their average (dotted black) for three training strategies, from left to right: no-replay, interleaved, and pseudo-optimal. The rightmost panel shows the average loss over the entire training.
@@
Fig. 6 shows the training loss under the different training protocols for $\gamma=0.5$. While the no-replay strategy appears to be successful for small downstream datasets (i.e., a few epochs in the online framework) in the longer run it leads to strong forgetting and high average loss. The interleaved is beneficial in the long run but largely slows down learning of the new task. Overall, the pseudo-optimal protocol identified in Sec. 3.1 shows a better performance over the entire trajectory.
@@
This result is not limited to the specific value of $\gamma$. In Fig. 7 we show snapshots of the average loss for different downstream task sizes while spanning over a range of $\gamma$s. This figure provides additional support to the observation reported previously that the no-replay strategy is optimal for small downstream tasks, the interleaved strategy is convenient for large downstream tasks, and the pseudo-optimal one combines the benefits of the two leading to the best performance overall. In summary, the pseudo-optimal strategy derived for the synthetic model performs well on real-world data. Notably, despite the differences between the synthetic and real settings—such as data structure—the pseudo-optimal strategy remains effective, demonstrating its robustness across problems.
@@ Conclusion.
This work marks an initial step toward a systematic theoretical understanding of optimal training protocols in synthetic learning settings. Here, we discuss limitations and highlight promising future directions. First, Pontryagin’s maximum principle provides a necessary condition for optimality but does not guarantee a global optimum. Nevertheless, the strategies derived from this approach performed significantly better than previously proposed heuristics. Additionally, Pontryagin’s principle does not easily extend to stochastic problems. This limitation is overcome in the high-dimensional limit where concentration results provide deterministic dynamical equations. For simplicity, we focus on i.i.d. Gaussian inputs, but our analysis can be extended to more structured data models (Goldt et al., 2020; Loureiro et al., 2021; Adomaityte et al., 2023) to study how input distribution affects task selection. In particular, we do not model the relative task difficulty—an important extension that naturally connects to the theory of curriculum learning (Weinshall & Amir, 2020; Saglietti et al., 2022; Cornacchia & Mossel, 2023; Abbe et al., 2023). Furthermore, it would be interesting to go beyond the study of online dynamics to understand the impact of memorisation in batch learning settings (Sagawa et al., 2020). Existing results in the spurious correlations (Ye et al., 2021) and fairness (Ganesh et al., 2023) literature suggest a strong dependence of the classifier’s bias on the presentation order in batch learning. Our method can be applied to mean-field models—like (Mannelli et al., 2024; Jain et al., 2024)—to theoretically investigate this phenomenon. An interesting extension of our work involves applying recently-developed statistical physics methods to the study of deeper networks (Bordelon & Pehlevan, 2022b; a) and more complex learning architectures (Rende et al., 2024; Tiberi et al., 2024). Another interesting direction concerns finding optimal protocols for shaping, where task order significantly impacts both animal learning and neural networks (Skinner, 1938; Tong et al., 2023; Lee et al., 2024).
@@ Appendix A Details on the theoretical derivations
Notice that, due to the online-learning setup, at each training epoch the input $\bm{x}$ is independent of the weights. Therefore, due to the Gaussianity of the inputs, the local fields are also jointly Gaussian with zero mean and second moments given by the overlaps:

### v2

@@ 3.1 Experiments on synthetic data
We use the equations derived in Sec. 2 to study optimal replay during the second phase of training. To this end, we take the task-selection variable as our control $u(\alpha)\!=\!t_{c}(\alpha)\!\in\!\{1,2\}$, while we set $t_{c}\!=\!1$ during the first training phase. The result of the optimisation in Eq. 10 strikes the balance between training on the new task and replaying the old task. We do not enforce any constraints on the number of samples from task $1$ to use in the second phase. Therefore, our method provides both the optimal fraction of replayed samples and the optimal task ordering, depending on the time window $\alpha_{F}$. Fig. 2 compares the learning dynamics of three different strategies, depicting the loss on task $1$ (full orange line), task $2$ (dashed green line), and their average (dotted black line) as a function of the training time $\alpha$. Numerical simulations—marked by cross symbols—are in excellent agreement with our theory. Deviations are smaller than $1/\sqrt{N}$, compatible with finite-size effects.
@@
To understand the structure of the optimal strategy, we examine its performance in relation to task similarity $\gamma$. Fig. 3a) depicts the average loss at the end of training as a function of $\gamma$. For the no-replay strategy, we reproduce the findings from previous works (Lee et al., 2021; 2022): the highest error occurs at intermediate task similarity. Lee et al. (2022) explained this non-monotonicity as a trade-off between node re-use and node activation. Specifically, for small $\gamma$, there is minimal interference between tasks. One hidden neuron predominantly aligns with the new task, while the other neuron retains the knowledge of the old task, leading to task specialisation. At large $\gamma$, features from task $1$ are reused for task $2$, avoiding forgetting. However, at intermediate $\gamma$, interference is maximal: both neurons quickly align with task $2$, and task $1$ is forgotten. Remarkably, Fig. 3a) shows that replay reverses this trend, with the minimal error occurring at intermediate $\gamma$. To explain this nontrivial behaviour, we must first understand the optimal replay protocol.
@@
The optimal replay dynamics is illustrated in panels b-e) of Fig. 3 and displays a highly structured protocol. We can interpret this structure a posteriori: an initial focus phase without replay is followed by a revision phase involving interleaved replay. The transition between these two phases corresponds approximately to the point at which the loss on the new task matches the loss on the old one. To investigate the significance of this structure, we also test an interleaved strategy, plotted in Fig. 3a). In this case, the task ordering in the second training phase is fully randomised while maintaining the same overall replay fraction of the optimal strategy. This protocol has a performance gap compared to the optimal one, showing the importance of a properly structured replay scheme. Additionally, we test a “pseudo-optimal” variant, where the focus phase is retained, but the revision phase is randomised. This variant performs comparably to the optimal strategy, suggesting that while the specific order of the revision phase is largely unimportant, it is key to precede it with a training phase on the new task.
@@
We can now attempt to understand the inverted non-monotonic behaviour of the average loss as a function of $\gamma$ under the optimal protocol. First, as shown in Fig. 8 of Appendix C, the optimal protocol achieves a good level of node specialisation across all values of $\gamma$. Thus, replay prevents the task interference that typically causes performance deterioration at intermediate $\gamma$. The non-monotonic behaviour of the optimal curve in Fig. 3a) arises from a different origin, involving two opposing effects related to the first-layer weights and the readout. The initial decrease of the loss with $\gamma$ is quite intuitive, as only minimal knowledge can be transferred from task $1$ to task $2$ when $\gamma$ is small. Consequently, the focus phase on task $2$ must be longer for smaller $\gamma$, leaving less time to revise task $1$, thereby reducing performance. On the other hand, the performance decrease observed in Fig. 3a) for $\gamma>0.3$ is more subtle and is related to the readout layer. Once the two hidden neurons have specialised—each aligning with one of the teacher vectors—we expect the readout weights corresponding to the incorrect teacher to be suppressed. Specifically, if $\bm{w}_{1}=\bm{w}_{*}^{(1)}$ and $\bm{w}_{2}=\bm{w}_{*}^{(2)}$, the learning dynamics should drive the readout weights $\bm{v}^{(1)}=(v^{(1)}_{1},v^{(1)}_{2})^{\top}$ and $\bm{v}^{(2)}=(v^{(2)}_{1},v^{(2)}_{2})^{\top}$ towards $\bm{v}^{(1)}=(1,0)^{\top}$ and $\bm{v}^{(2)}=(0,1)^{\top}$ to achieve full recovery of the teacher networks. As shown in Fig. 8 of Appendix C, the time required to suppress the off-diagonal weights $v^{(1)}_{2}$ and $v^{(2)}_{1}$ increases as $\gamma\to 1$. This is intuitive, as higher task similarity $\gamma$ reduces the distinction between tasks, slowing the suppression of the off-diagonal weights. In Appendix B, we derive analytically the convergence timescale $\alpha_{\rm conv}$ of the readout layer, showing that
@@
While the above results are obtained with a constant learning rate, annealing schedules are widely used in machine learning. Thus, it is relevant to study the optimal interplay between replay and learning rate. Optimal learning rate dynamics have been studied with a similar approach in Saad & Rattray (1997). We jointly optimise the task-selection protocol together with the learning rate to investigate its impact on continual learning. Fig. 4 shows the optimal learning rate schedule for task similarity $\gamma\!=\!0.3$ in the second training phase of duration $\alpha_{F}\!=\!25$. Similarly as for constant learning rate (panel c) of Fig. 3), optimal task-selection is characterised by an initial focus phase. Notably, this phase coincides with a strong annealing of the learning rate to achieve optimal performance. Intuitively, when learning the new task, the learning rate starts high and is gradually decreased over time. Interestingly, while entering the revision phase, the optimal learning rate schedule exhibits a highly nontrivial structure (see Fig. 4). Indeed, although the optimal learning rate curve is unique, we find that effectively it can be seen as two different curves, associated to the respective tasks. In practice, the optimal learning rate curve “jumps” between these two curves according to the task selected at a given training time. performance. Fig. 9 of Appendix C shows that the joint optimization of learning rate and task selection outperforms all other protocols, including exponential and power-law learning rate schedules combined with interleaved replay.
@@
We also consider a multi-task setting (see Fig. 5) where both tasks must be learned from scratch within a fixed number of steps, corresponding to a total training time $\alpha_{F}$. We first consider sequential learning, i.e., training only on task $1$ for $\alpha<\alpha_{F}/2$, then only on task $2$, or vice versa. As shown in Fig. 5a) sequential learning leads to catastrophic forgetting, with the worst performance observed at intermediate task similarity. In contrast, a randomly interleaved strategy, where examples from both tasks are presented in equal proportion but in random order, shows significant improvement. This approach can exploit task similarity, leading to a monotonic decrease in average loss as $\gamma$ increases. The optimal strategy, displayed in Fig. 5b-e) for various values of $\gamma$, follows a structured interleaved protocol that further enhances performance. Contrary to the continual learning framework, the optimal structure gives only marginal gain over the plain interleaved strategy. This observation aligns with our pseudo-optimal strategy, which suggests employing interleaved replay once performance on both tasks becomes comparable.
@@ 3.2 Experiments on real data
We consider the experimental framework established in Ramasesh et al. (2020); Lee et al. (2022) for the study of task similarity in relation to catastrophic forgetting. We use the Fashion-MNIST dataset (Xiao et al., 2017) to generate upstream and downstream tasks. The upstream dataset—$\mathcal{D}_{1}=\{\boldsymbol{x}_{i}^{(1)},y_{i}^{(1)}\}_{i}$—consists in a pair of classes from the standard dataset. The downstream dataset is generated by a linear interpolation of the upstream dataset with a second auxiliary dataset—$\tilde{\mathcal{D}}=\{\tilde{\boldsymbol{x}}_{i},\tilde{y}_{i}\}_{i}$—containing a new pair of classes,
$\mathcal{D}_{2}=\{\boldsymbol{x}_{i}^{(2)},y_{i}^{(2)}\}_{i}=\{\gamma\boldsymbol{x}_{i}^{(1)}+(1-\gamma)\tilde{\boldsymbol{x}}_{i},\gamma y_{i}^{(1)}+(1-\gamma)\tilde{y}_{i}\}_{i},$ (12)
@@
where the parameter $\gamma$ controls the task similarity. We then train a standard two-layer feedforward ReLU neural network on the two datasets using online SGD on a squared error loss. We consider a dynamical multi-head architecture (Zhou et al., 2012; Rusu et al., 2016) where the readout weights are changed switching from one task to another, but the hidden layer is shared. During training, we apply the three strategies discussed in the previous sections: a no-replay strategy, a strategy with interleaved replay, and a “pseudo-optimal” strategy. Recall that the latter is inspired by the optimal protocol derived in the previous section. It consists of an initial phase of training exclusively on the new task until performance on both tasks becomes comparable, followed by a phase of interleaved replay. Crucially, this protocol can be easily implemented in practice, as it only requires an estimate of the generalisation error on the two tasks, which can be obtained in real-world settings.
@@
Figure 6: Training dynamics. Training curves on the modified fashion MNIST task at similarity $\gamma=0.5$. The network is trained for $10.000$ steps on the first task before switching to the second task and being trained for additional $10.000$ steps. The results are obtained from $100$ realisations of the problem. The first three panels show the test loss on task $1$ (solid orange), task $2$ (dashed green), and their average (dotted black) for three training strategies, from left to right: no-replay, interleaved, and pseudo-optimal. The rightmost panel shows the average loss over the entire training.
@@
Fig. 6 shows the training loss under the different training protocols for $\gamma=0.5$. While the no-replay strategy appears to be successful for small downstream datasets (i.e., a few training steps in the online framework) in the longer run it leads to strong forgetting and high average loss. This behavior is intuitive: for small datasets, the initial loss on the new task is high, leading to a substantial decrease in loss early on, which temporarily outweighs the decline in performance on the previous task. The interleaved is beneficial in the long run but largely slows down learning of the new task. Overall, the pseudo-optimal protocol identified in Sec. 3.1 shows a better performance over the entire trajectory.
@@
This result is not limited to the specific value of $\gamma$. Fig. 7 shows the average final loss as a function of $\gamma$. From left to right, different panels correspond to an increasing number of available samples for the downstream task, comparable to $\alpha_{F}$ in the theoretical model. For small downstream tasks, the no-replay strategy is optimal, as shown in the second and third panels. As the size of the downstream task increases, the interleaved strategy approaches optimal performance while the no-replay becomes suboptimal. The pseudo-optimal strategy combines the advantages of both approaches, interpolating between no-replay and interleaved strategies to automatically adapt to the computational budget. This results in the best overall performance across regimes. We confirm this observation with additional experiments on CIFAR10 in Appendix C.2. Notably, despite the differences between the synthetic and real settings—such as data structure—the pseudo-optimal strategy remains effective on real-world data, demonstrating its robustness and broad applicability.
@@ Conclusion.
This work takes a first step toward understanding the theory behind optimal training protocols for neural networks. In the following, we discuss current limitations and outline promising directions for future research. First, Pontryagin’s maximum principle provides a necessary condition for optimality but does not guarantee a global optimum. Nevertheless, the strategies derived from this approach in the settings under consideration perform significantly better than previously proposed heuristics. Additionally, Pontryagin’s principle does not easily extend to stochastic problems. This limitation is overcome in the high-dimensional limit where concentration results provide deterministic dynamical equations. For simplicity, we focus on i.i.d. Gaussian inputs, but our analysis can be extended to more structured data models (Goldt et al., 2020; Loureiro et al., 2021; Adomaityte et al., 2023) to study how input distribution affects task selection. In particular, we do not model the relative task difficulty—an important extension that naturally connects to the theory of curriculum learning (Weinshall & Amir, 2020; Saglietti et al., 2022; Cornacchia & Mossel, 2023; Abbe et al., 2023). Furthermore, it would be interesting to go beyond the study of online dynamics to understand the impact of memorisation in batch learning settings (Sagawa et al., 2020). Existing results in the spurious correlations (Ye et al., 2021) and fairness (Ganesh et al., 2023) literature suggest a strong dependence of the classifier’s bias on the presentation order in batch learning. Our method can be applied to mean-field models—like (Mannelli et al., 2024; Jain et al., 2024)—to theoretically investigate this phenomenon. An interesting extension of our work involves applying recently-developed statistical physics methods to the study of deeper networks and more complex learning architectures (Bordelon & Pehlevan, 2022b; a; Rende et al., 2024; Tiberi et al., 2024). Another interesting direction concerns finding optimal protocols for shaping, where task order significantly impacts both animal learning and neural networks (Skinner, 1938; Tong et al., 2023; Lee et al., 2024).
@@ Appendix A Details on the theoretical derivations
In the main text, we focus on the case $K=T$, where the student has, in principle, the capacity to perfectly solve the problem and represent all teachers. Specifically, there exists a configuration of the student’s parameters that achieves perfect recovery. This configuration corresponds to aligning each of the student’s hidden neurons with a specific teacher/task. Explicitly, this configuration is given by $\bm{w}_{k}=\bm{w}_{*}^{(k)}$ and $v_{k}^{(t)}=\delta_{k,t}$, where $\delta_{k,t}$ denotes the Kronecker delta. However, our theory remains valid for arbitrary $K$ and $T$.
@@
Notice that, due to the online-learning setup, at each training step the input $\bm{x}$ is independent of the weights. Therefore, due to the Gaussianity of the inputs, the local fields are also jointly Gaussian with zero mean and second moments given by the overlaps: